OpenStudy (anonymous):

3(a + b) = 3a + 3b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (anonymous):

The distributive property.

OpenStudy (anonymous):

Distributive: 3(a + b) = 3a + 3b Commutative (Addition): a + b + c = c + a + b Associative (Addition): (a + b) + c = a + (b + c) Symmetric (Addition): a = b, b = a